Tom buys a DVD player on 12 monthly payments. The cash price is $600. He pays 10% deposit and the store charges 9% interest
1.) what is the monthly payments?
1
Expert's answer
2018-09-12T09:40:08-0400
Tom buys a DVD player on 12 monthly payments. The cash price is $600.
If he pays 10% deposit ($60) and the store charges 9% interest, then the monthly payments are...
Future value or loan value is FV = PV*(1 + i)^t,
where PV is present value or price, i is rate of interest, t is the number of years.
Monthly payment is MP = FV/n,
where n is the number of payments;
MP = (600 - 60)*(1+0.09)/12 = $49.05.